OpenStudy (anonymous):

Complete the ordered pair below for the equation y=-3x+9 (0,_) _____________________________________________________________ Complete the ordered pair. (0,_) (Type an integer or a simplified fraction.

OpenStudy (anonymous):

y=-3(0)+9 y=0+9 y=9

OpenStudy (anonymous):

(0,9)